sorry i put this in earlier i 4got 2 add sumthin
8t-r=12t, for t

Answer by tutorcecilia(2152)

8t-r=12t
8t-8t-r=12t-8t
-r=4t
-r/-1=4t/-1
r=-4t
.
Check by plugging (r=-4t) back into the equation.